# Prox-DBRO-VR: A Unified Analysis on Decentralized Byzantine-Resilient Composite Stochastic Optimization with Variance Reduction and Non-Asymptotic Convergence Rates

### v6

@@ VI Conclusions
$\left.{\left.{\left.{-\alpha{\partial_{x}}\chi\left({{x^{*}}}\right)}\right\}}\right\|_{2}^{2}}\right]$
@@
$+\left.{\left.{{\partial_{x}}\delta\left({{x_{k}}}\right)}\right\|_{2}^{2}}\right],$
where the inequality applies the relationship (12). We continue to seek an upper bound for ${\mathbb{E}_{k}}\left[{\left\|{r_{k}-\nabla F\left({{x^{*}}}\right)+{\partial_{x}}\chi\left({{x_{k}}}\right)-{\partial_{x}}\chi\left({{x^{*}}}\right)+{\partial_{x}}\delta\left({{x_{k}}}\right)}\right\|_{2}^{2}}\right]$ as follows:
@@
where the first inequality applies ${\left\|{c+d}\right\|^{2}}\leq 2{c^{2}}+2{d^{2}}$ twice, and the second equality employs Lemma 2. To proceed, we bound $\left\|{{\partial_{x}}\delta\left({{x_{k}}}\right)}\right\|_{2}^{2}$ as follows:
$\left\|{{\partial_{x}}\delta\left({{x_{k}}}\right)}\right\|_{2}^{2}=$ $\sum\limits_{i\in\mathcal{R}}{\left\|{{\phi_{i}}\sum\limits_{j\in{\mathcal{B}_{i}}}{{\partial_{{x_{i}}}}{{\left\|{{x_{i,k}}-{z_{ij,k}}}\right\|}_{a}}}}\right\|}_{2}^{2}$ (47)
$\leq$ $n{{\bar{\phi}}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}},$
where the inequality holds true, since the $b$-norm ($b\geq 1$) of ${\partial_{{x_{i}}}}{\left\|{{x_{i,k}}-{z_{ij,k}}}\right\|_{a}}$, $\forall i\in\mathcal{R}$, is no larger than $1$ owing to Proposition 1, i.e.,
@@
Following the same technical line of (47)-(48), it is not difficult to verify
$\left\|{{\partial_{x}}\chi\left({{x_{k}}}\right)-{\partial_{x}}\chi\left({{x^{*}}}\right)}\right\|_{2}^{2}$
$=$ $\sum\limits_{i\in\mathcal{R}}{\left\|{{\phi_{i}}\sum\limits_{j\in{\mathcal{R}_{i}}}{\left({{\partial_{{x_{i}}}}{{\left\|{{x_{i,k}}-{x_{j,k}}}\right\|}_{a}}-{\partial_{{x_{i}}}}{{\left\|{x_{i}^{*}-x_{j}^{*}}\right\|}_{a}}}\right)}}\right\|_{2}^{2}}$
$\leq$ $4n{{\bar{\phi}}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}.$ (49)
@@
$\leq$ $4\left({2Lt_{k}+\left({2L-\mu}\right){D_{F}}\left({{x_{k}},{x^{*}}}\right)}\right)+8n{{\bar{\phi}}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}$
$+4n{{\bar{\phi}}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}.$
Since the local objective function ${f_{i}}\left({{x_{i}}}\right)$, $\forall i\in\mathcal{R}$, is $\mu$-strongly convex and $L$-smooth according to Assumption 1, we have
@@
$\leq$ $\frac{{\mu L}}{{\mu+L}}\left\|{{x_{k}}-{x^{*}}}\right\|_{2}^{2}+\frac{{1}}{{\mu+L}}\left\|{\nabla F\left({{x_{k}}}\right)-\nabla F\left({{x^{*}}}\right)}\right\|_{2}^{2}.$
Recalling the definition of $\chi\left({{x_{k}}}\right)$, we know that $\chi\left({{x_{k}}}\right)$ is a convex function. Therefore, it is straightforward to obtain
@@
We next analyze the term $-2\left\langle{{x_{k}}-{x^{*}},\partial\delta\left({{x_{k}}}\right)}\right\rangle$ with the aid of an arbitrary constant $\gamma>0$,
$-2\left\langle{{x_{k}}-{x^{*}},\partial\delta\left({{x_{k}}}\right)}\right\rangle\leq\gamma\left\|{{x_{k}}-{x^{*}}}\right\|_{2}^{2}+\frac{{n{{\bar{\phi}}^{2}}}}{\gamma}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}},$ (53)
where we apply the Young’s inequality and (47). Plugging the results (50)-(53) into (45) gives
@@
$\leq$ $\left({1-\left({\frac{{2\mu L}}{{\mu+L}}-\gamma}\right)\alpha}\right)\left\|{{x_{k}}-{x^{*}}}\right\|_{2}^{2}+8n{{\bar{\phi}}^{2}}{\alpha^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}$
$+4{\alpha^{2}}\left({2Lt_{k}+\left({2L-\mu}\right){D_{F}}\left({{x_{k}},{x^{*}}}\right)}\right)+4n{{\bar{\phi}}^{2}}{\alpha^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}$
$+\frac{{n\alpha}}{\gamma}{{\bar{\phi}}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}.$
Via setting $\gamma=\mu L/\left({\mu+L}\right)$, we can rewrite (54) as follows:
@@
$+8n{{\bar{\phi}}^{2}}{\alpha^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}+4n{{\bar{\phi}}^{2}}{\alpha^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}\!+\!\frac{{n\alpha}}{\gamma}{{\bar{\phi}}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}.$
@@
${\mathbb{E}_{k}}\left[{\left\|{{x_{k+1}}-{x^{*}}}\right\|_{2}^{2}}\right]+c\left({{\mathbb{E}_{k}}\left[{t_{k+1}}\right]-t_{k}}\right)$
$\leq$ $\left({1-\gamma{\alpha}}\right)\left\|{{x_{k}}-{x^{*}}}\right\|_{2}^{2}+4n{{\bar{\phi}}^{2}}\alpha^{2}({2\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}+\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}})$
$+\frac{{n{{\bar{\phi}}^{2}}}}{\gamma}{\alpha}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}+8L\alpha^{2}t_{k}+4\left({2L-\mu}\right){\alpha^{2}}{D_{F}}\left({{x_{k}},{x^{*}}}\right)$ (57)
@@
where the last inequality employs the $L$-smoothness of the local objective function $f_{i}$, $\forall i\in\mathcal{R}$. We proceed by choosing $0<\alpha\leq\gamma/\left({8L\left({2L-\mu}\right)}\right)$ and setting ${{c}=\tilde{c}{\alpha}}$ with $0<\tilde{c}\leq{q_{\min}}\gamma/L$, such that (57) becomes
@@
$\leq$ $\left({1-\frac{\gamma}{4}{\alpha}}\right)\left\|{{x_{k}}-{x^{*}}}\right\|_{2}^{2}+\left({8L{\alpha^{2}}-\frac{c}{{{q_{\max}}}}}\right)t_{k}+P_{1}^{c}{\alpha^{2}}$
$+P_{2}\alpha.$
To proceed, via fixing $\tilde{c}={q_{\min}}\gamma/L$ and $0<\alpha\leq\gamma/\left({8L\left({2L-\mu}\right)}\right)$, (58) is equivalent to
@@
$\leq$ $\left({1-\frac{\gamma}{4}{\alpha}}\right)\left\|{{x_{k}}-{x^{*}}}\right\|_{2}^{2}+4n{{\bar{\phi}}^{2}}\alpha^{2}({2\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}+\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}})$
$+\frac{{n{{\bar{\phi}}^{2}}}}{\gamma}{\alpha}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}+\left({8L{\alpha}-\frac{{\gamma{q_{\min}}}}{{L{q_{\max}}}}}\right){\alpha}t_{k}.$
@@
${\mathbb{E}_{k}}\left[{{U_{k+1}}}\right]$ (60)
$\leq$ ${\mathbb{E}_{k}}\left[{\left\|{{x_{k+1}}-{x^{*}}}\right\|_{2}^{2}}\right]+\frac{{{q_{\min}}\gamma\alpha}}{L}{\mathbb{E}_{k}}\left[{t_{k+1}}\right]$
$\leq$ $\left({1-\frac{\gamma}{4}\alpha}\right)\left\|{{x_{k}}-{x^{*}}}\right\|_{2}^{2}+\left({1-\frac{\gamma}{4}\alpha}\right)\frac{{{q_{\min}}\gamma\alpha}}{L}t_{k}+P_{1}^{c}{\alpha^{2}}$
$+{P_{2}}\alpha$
@@
Summarizing all the upper bounds on the constant step-size generates a feasible selection range as follows:
@@
Based on (61), taking the full expectation on the both sides of (60) obtains
@@
For $\forall k\geq 0$, applying telescopic cancellation to (62) obtains
${\mathbb{E}}\left[{\left\|{{x_{k+1}}-{x^{*}}}\right\|_{2}^{2}}\right]$ (63)
$\leq$ ${\left({1-\frac{\gamma}{4}\alpha}\right)^{k+1}}{U_{0}}+4\left({\frac{{{P_{1}}}}{\gamma}\alpha+E}\right)\left({1-{{\left({1-\frac{\gamma}{4}\alpha}\right)}^{k+1}}}\right),$
@@
In view of the compact form (IV) associated with the proposed algorithms, we make a transformation as follows:
${x_{k+1}}=$ $\arg\mathop{\min}\limits_{y\in{\mathbb{R}^{\left|\mathcal{R}\right|n}}}\left\{{G\left(y\right)+\frac{1}{{2{\alpha_{k}}}}\left\|{y-{x_{k}}+{\alpha_{k}}\left({{r_{k}}}\right.}\right.}\right.$
@@
$+{\alpha_{k}}{\partial_{x}}G\left({{x_{k+1}}}\right).$
This implies that if ${x_{k+1}}$ is the minimizer of the next update of the proposed algorithm, we are guaranteed to obtain a vector ${G_{k+1}}\in{\partial_{x}}G\left({{x_{k+1}}}\right)$, such that
${x_{k+1}}-{x_{k}}+{\alpha_{k}}\left({{r_{k}}+{G_{k+1}}+{\partial_{x}}\chi\left({{x_{k}}}\right)+{\partial_{x}}\delta\left({{x_{k}}}\right)}\right)={0_{\left|\mathcal{R}\right|n}},$
@@
${x_{k+1}}={x_{k}}-{\alpha_{k}}\left({{r_{k}}+{G_{k+1}}+{\partial_{x}}\chi\left({{x_{k}}}\right)+{\partial_{x}}\delta\left({{x_{k}}}\right)}\right).$ (64)
We next proceed with the convergence analysis based on the transformed version (64) of the compact form (IV).
@@
$=$ $\left\|{{x_{k}}-{x^{*}}}\right\|_{2}^{2}-2{\alpha_{k}}{\mathbb{E}_{k}}\left[{\left\langle{{x_{k}}-{x^{*}},{r_{k}}+{\partial_{x}}\chi\left({{x_{k}}}\right)+{G_{k+1}}}\right\rangle}\right]$
$+\alpha_{k}^{2}{\mathbb{E}_{k}}\left[{\left\|{{r_{k}}+{G_{k+1}}+{\partial_{x}}\chi\left({{x_{k}}}\right)+{\partial_{x}}\delta\left({{x_{k}}}\right)}\right\|_{2}^{2}}\right]$
@@
Considering ${G^{*}}\in{\partial_{x}}G\left({{x^{*}}}\right)$ and the optimality condition $\nabla F\left({{x^{*}}}\right)+{G^{*}}+{\partial_{x}}\chi\left({{x^{*}}}\right)={0_{mn}}$, we continue to seek an upper bound for ${\mathbb{E}_{k}}\left[{\left\|{{r_{k}}+{G_{k+1}}+{\partial_{x}}\chi\left({{x_{k}}}\right)+{\partial_{x}}\delta\left({{x_{k}}}\right)}\right\|_{2}^{2}}\right]$ as follows:
${\mathbb{E}_{k}}\left[{\left\|{{r_{k}}+{G_{k+1}}+{\partial_{x}}\chi\left({{x_{k}}}\right)+{\partial_{x}}\delta\left({{x_{k}}}\right)}\right\|_{2}^{2}}\right]$ (66)
$=$ ${\mathbb{E}_{k}}\left[{\left\|{{r_{k}}-\nabla F\left({{x^{*}}}\right)+{G_{k+1}}-{G^{*}}+{\partial_{x}}\chi\left({{x_{k}}}\right)-{\partial_{x}}\chi\left({{x^{*}}}\right)}\right.}\right.$
$\left.+{\left.{{\partial_{x}}\delta\left({{x_{k}}}\right)}\right\|_{2}^{2}}\right]$
$\leq$ $4{\mathbb{E}_{k}}\left[{\left\|{{r_{k}}-\nabla F\left({{x^{*}}}\right)}\right\|_{2}^{2}}\right]+4\left\|{{G_{k+1}}-{G^{*}}}\right\|_{2}^{2}$
@@
$\leq$ $4{\mathbb{E}_{k}}\left[{\left\|{{r_{k}}-\nabla F\left({{x^{*}}}\right)}\right\|_{2}^{2}}\right]+4\left\|{{G_{k+1}}-{G^{*}}}\right\|_{2}^{2}$
$+16n{{\bar{\phi}}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}+4n{{\bar{\phi}}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}$
@@
$+16n{{\bar{\phi}}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}+4n{{\bar{\phi}}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}},$
where the second inequality uses the results (47) and (49), and the last inequality is owing to Lemma 2 and Assumption 3. To proceed, recalling the definition of ${{\partial_{x}}G\left({{x_{k}}}\right)}$, it is not difficult to verify
$-\left\langle{{x_{k}}-{x^{*}},{G_{k+1}}-{G^{*}}}\right\rangle\leq 0,\$ (67)
which is owing to the convexity of ${g}\left({\tilde{x}}\right)$, $\forall i\in\mathcal{R}$. Based on the relationships (52) and (67), we know that
$-2{\mathbb{E}_{k}}\left[{\left\langle{{x_{k}}-{x^{*}},{r_{k}}+{G_{k+1}}+{\partial_{x}}\chi\left({{x_{k}}}\right)}\right\rangle}\right]$ (68)
@@
$-2{\mathbb{E}_{k}}\left[{\left\langle{{x_{k}}-{x^{*}},{G_{k+1}}-{G^{*}}+{\partial_{x}}\chi\left({{x_{k}}}\right)-{\partial_{x}}\chi\left({{x^{*}}}\right)}\right\rangle}\right]$

### v12

@@ VI Conclusion
$-\left.{\left.{\left.{\alpha{\partial_{x}}\chi\left({{x^{*}}}\right)}\right\}}\right\|_{2}^{2}}\right]$
@@
$+\left.{\left.{{\partial_{x}}\delta\left({{x_{k}}}\right)}\right\|_{2}^{2}}\right],$ (46)
where the second inequality applies (12). We continue to seek an upper bound for ${\mathbb{E}_{k}}\left[{\left\|{r_{k}-\nabla F\left({{x^{*}}}\right)+{\partial_{x}}\chi\left({{x_{k}}}\right)-{\partial_{x}}\chi\left({{x^{*}}}\right)+{\partial_{x}}\delta\left({{x_{k}}}\right)}\right\|_{2}^{2}}\right]$ in the sequel.
@@
where the first inequality applies ${\left\|{c+d}\right\|^{2}}\leq 2{c^{2}}+2{d^{2}}$ twice, and the second equality uses Lemma 2. To proceed, we bound $\left\|{{\partial_{x}}\delta\left({{x_{k}}}\right)}\right\|_{2}^{2}$ as follows:
$\left\|{{\partial_{x}}\delta\left({{x_{k}}}\right)}\right\|_{2}^{2}=$ $\sum\limits_{i\in\mathcal{R}}{\left\|{{\phi}\sum\limits_{j\in{\mathcal{B}_{i}}}{{\partial_{{x_{i}}}}{{\left\|{{x_{i,k}}-{z_{ij,k}}}\right\|}_{a}}}}\right\|}_{2}^{2}$ (48)
$\leq$ $n{{\phi}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}},$
where the inequality holds true since the $b$-norm ($b\geq 1$) of ${\partial_{{x_{i}}}}{\left\|{{x_{i,k}}-{z_{ij,k}}}\right\|_{a}}$, $\forall i\in\mathcal{R}$, is no larger than $1$ owing to Proposition 1, i.e.,
@@
Following the same technical line as in (48)-(49), it is not difficult to verify
$\left\|{{\partial_{x}}\chi\left({{x_{k}}}\right)-{\partial_{x}}\chi\left({{x^{*}}}\right)}\right\|_{2}^{2}$ (50)
$=$ $\sum\limits_{i\in\mathcal{R}}{\left\|{{\phi}\sum\limits_{j\in{\mathcal{R}_{i}}}{\left({{\partial_{{x_{i}}}}{{\left\|{{x_{i,k}}-{x_{j,k}}}\right\|}_{a}}-{\partial_{{x_{i}}}}{{\left\|{x_{i}^{*}-x_{j}^{*}}\right\|}_{a}}}\right)}}\right\|_{2}^{2}}$
$\leq$ $4n{{\phi}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}.$
@@
$\leq$ $4\left({2Lt_{k}+\left({2L-\mu}\right){D_{F}}\left({{x_{k}},{x^{*}}}\right)}\right)+8n{{\phi}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}$
$+4n{{\phi}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}.$
Since the local objective functions ${f_{i}}\left({{x_{i}}}\right)$, $\forall i\in\mathcal{R}$, are $\mu$-strongly convex and $L$-smooth according to Assumption 1, we have
@@
$\leq$ $-\frac{{\mu L}}{{\mu+L}}\left\|{{x_{k}}-{x^{*}}}\right\|_{2}^{2}-\frac{{1}}{{\mu+L}}\left\|{\nabla F\left({{x_{k}}}\right)-\nabla F\left({{x^{*}}}\right)}\right\|_{2}^{2}.$
Recalling the definition of $\chi\left({{x_{k}}}\right)$, we know that it is a convex function. Therefore, it is straightforward to obtain
@@
Applying the Young’s inequality and the relation (48) to the term $-2\left\langle{{x_{k}}-{x^{*}},\partial\delta\left({{x_{k}}}\right)}\right\rangle$ yields
$-2\left\langle{{x_{k}}-{x^{*}},\partial\delta\left({{x_{k}}}\right)}\right\rangle\leq\gamma\left\|{{x_{k}}-{x^{*}}}\right\|_{2}^{2}+\frac{{n{{\phi}^{2}}}}{\gamma}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}},$ (54)
where the constant $\gamma>0$. Plugging the results (51)-(54) into (46) gives
@@
$\leq$ $\left({1-\left({\frac{{2\mu L}}{{\mu+L}}-\gamma}\right)\alpha}\right)\left\|{{x_{k}}-{x^{*}}}\right\|_{2}^{2}+8n{{\phi}^{2}}{\alpha^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}$
$+4{\alpha^{2}}\left({2Lt_{k}+\left({2L-\mu}\right){D_{F}}\left({{x_{k}},{x^{*}}}\right)}\right)+4n{{\phi}^{2}}{\alpha^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}$
$+\frac{{n\alpha}}{\gamma}{{\phi}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}.$
By setting $\gamma=\mu L/\left({\mu+L}\right)$, we can rewrite (55) as follows:
@@
$+8n{{\phi}^{2}}{\alpha^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}+4n{{\phi}^{2}}{\alpha^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}\!+\!\frac{{n\alpha}}{\gamma}{{\phi}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}.$
@@
${\mathbb{E}_{k}}\left[{\left\|{{x_{k+1}}-{x^{*}}}\right\|_{2}^{2}}\right]+c\left({{\mathbb{E}_{k}}\left[{t_{k+1}}\right]-t_{k}}\right)$ (58)
$\leq$ $\left({1-\gamma{\alpha}}\right)\left\|{{x_{k}}-{x^{*}}}\right\|_{2}^{2}+4n{{\phi}^{2}}\alpha^{2}({2\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}+\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}})$
$+\frac{{n{{\phi}^{2}}}}{\gamma}{\alpha}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}+8L\alpha^{2}t_{k}+4\left({2L-\mu}\right){\alpha^{2}}{D_{F}}\left({{x_{k}},{x^{*}}}\right)$
@@
where the last inequality employs $L$-smoothness of the local objective function $f_{i}$, $\forall i\in\mathcal{R}$. We proceed to choose $0<\alpha\leq\gamma/\left({8L\left({2L-\mu}\right)}\right)$ and set ${{c}=\tilde{c}{\alpha}}$ with $0<\tilde{c}\leq{q_{\min}}\gamma/L$, such that (58) becomes
@@
$\leq$ $\left({1\!-\!\frac{\gamma}{4}{\alpha}}\right)\left\|{{x_{k}}\!-\!{x^{*}}}\right\|_{2}^{2}\!+\!\left({8L{\alpha^{2}}\!-\!\frac{c}{{{q_{\max}}}}}\right)t_{k}+P_{1}^{c}{\alpha^{2}}\!+\!P_{2}\alpha.$
To proceed, via fixing $\tilde{c}={q_{\min}}\gamma/L$ and $0<\alpha\leq\gamma/\left({8L\left({2L-\mu}\right)}\right)$, it is equivalent to write (59) as
@@
$\leq$ $\left({1-\frac{\gamma}{4}{\alpha}}\right)\left\|{{x_{k}}-{x^{*}}}\right\|_{2}^{2}+4n{{\phi}^{2}}\alpha^{2}({2\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}+\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}})$
$+\frac{{n{{\phi}^{2}}}}{\gamma}{\alpha}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}+\left({8L{\alpha}-\frac{{\gamma{q_{\min}}}}{{L{q_{\max}}}}}\right){\alpha}t_{k}.$
@@
${\mathbb{E}_{k}}\left[{{U_{k+1}}}\right]\leq$ ${\mathbb{E}_{k}}\left[{\left\|{{x_{k+1}}-{x^{*}}}\right\|_{2}^{2}}\right]+\frac{{{q_{\min}}\gamma\alpha}}{L}{\mathbb{E}_{k}}\left[{t_{k+1}}\right]$ (61)
$\leq$ $\left({1-\frac{\gamma}{4}\alpha}\right)\left\|{{x_{k}}-{x^{*}}}\right\|_{2}^{2}\!+\!\left({1-\frac{\gamma}{4}\alpha}\right)\frac{{{q_{\min}}\gamma\alpha}}{L}t_{k}$
$+P_{1}^{c}{\alpha^{2}}+{P_{2}}\alpha$
@@
Summarizing all the upper bounds on the constant step-size generates a sufficient selection range as follows:
@@
Under the condition of (62), taking the total expectation on the both sides of (61) obtains
@@
Applying telescopic cancellation to (63) for $k\geq 0$ obtains
${\mathbb{E}}\left[{\left\|{{x_{k+1}}-{x^{*}}}\right\|_{2}^{2}}\right]\leq$ $4\left({\frac{{{P_{1}^{c}}}}{\gamma}\alpha+E}\right)\left({1-{{\left({1-\frac{\gamma}{4}\alpha}\right)}^{k+1}}}\right)$ (64)
$+{\left({1-\frac{\gamma}{4}\alpha}\right)^{k+1}}{U_{0}},$
@@
In view of the compact updates (7a)-(7b) of the proposed algorithms, we make a transformation as follows:
${x_{k+1}}=$ $\arg\mathop{\min}\limits_{y}\left\{{G\left(y\right)+\frac{1}{{2{\alpha_{k}}}}\left\|{y-{x_{k}}+{\alpha_{k}}\left({{r_{k}}}\right.}\right.}\right.$
@@
$+{\alpha_{k}}\partial G\left({{x_{k+1}}}\right).$
This implies that if ${x_{k+1}}$ is the minimizer of the next update, we are guaranteed to obtain a subgradient $G^{\prime}\left({{x_{k+1}}}\right)\in\partial G\left({{x_{k+1}}}\right)$ satisfying
${0_{\left|\mathcal{R}\right|n}}=$ ${x_{k+1}}-{x_{k}}+{\alpha_{k}}\left({{r_{k}}+{\partial_{x}}\chi\left({{x_{k}}}\right)+{\partial_{x}}\delta\left({{x_{k}}}\right)}\right)$
$+{\alpha_{k}}G^{\prime}\left({{x_{k+1}}}\right),$
@@
${x_{k+1}}={x_{k}}-{\alpha_{k}}\left({{r_{k}}+{G^{\prime}\left({{x_{k+1}}}\right)}+{\partial_{x}}\chi\left({{x_{k}}}\right)+{\partial_{x}}\delta\left({{x_{k}}}\right)}\right).$ (65)
We next proceed with the convergence analysis based on the transformed version (65) of the compact updates (7a)-(7b).
@@
$=$ $\left\|{{x_{k}}\!-\!{x^{*}}}\right\|_{2}^{2}\!-\!2{\alpha_{k}}{\mathbb{E}_{k}}\left[{\left\langle{{x_{k}}-{x^{*}},{r_{k}}\!+\!{\partial_{x}}\chi\left({{x_{k}}}\right)\!+\!{G^{\prime}\left({{x_{k+1}}}\right)}}\right\rangle}\right]$
$+\alpha_{k}^{2}{\mathbb{E}_{k}}\left[{\left\|{{r_{k}}+{G^{\prime}\left({{x_{k+1}}}\right)}+{\partial_{x}}\chi\left({{x_{k}}}\right)+{\partial_{x}}\delta\left({{x_{k}}}\right)}\right\|_{2}^{2}}\right]$
@@
Considering ${G^{\prime}\left({{x^{*}}}\right)}\in\partial G\left({{x^{*}}}\right)$ and the optimality condition $\nabla F\left({{x^{*}}}\right)+{G^{\prime}\left({{x^{*}}}\right)}+{\partial_{x}}\chi\left({{x^{*}}}\right)={0_{mn}}$, we continue to seek an upper bound for ${\mathbb{E}_{k}}\left[{\left\|{{r_{k}}+{G^{\prime}\left({{x_{k+1}}}\right)}+{\partial_{x}}\chi\left({{x_{k}}}\right)+{\partial_{x}}\delta\left({{x_{k}}}\right)}\right\|_{2}^{2}}\right]$ as follows:
${\mathbb{E}_{k}}\left[{\left\|{{r_{k}}+{G^{\prime}\left({{x_{k+1}}}\right)}+{\partial_{x}}\chi\left({{x_{k}}}\right)+{\partial_{x}}\delta\left({{x_{k}}}\right)}\right\|_{2}^{2}}\right]$ (67)
$=$ ${\mathbb{E}_{k}}\left[{\left\|{{r_{k}}-\nabla F\left({{x^{*}}}\right)+{G^{\prime}\left({{x_{k+1}}}\right)}-{G^{\prime}\left({{x^{*}}}\right)}+{\partial_{x}}\chi\left({{x_{k}}}\right)}\right.}\right.$
$\left.-{\partial_{x}}\chi\left({{x^{*}}}\right)+{\left.{{\partial_{x}}\delta\left({{x_{k}}}\right)}\right\|_{2}^{2}}\right]$
$\leq$ $4{\mathbb{E}_{k}}\left[{\left\|{{r_{k}}-\nabla F\left({{x^{*}}}\right)}\right\|_{2}^{2}}\right]+4\left\|{{G^{\prime}\left({{x_{k+1}}}\right)}-{G^{\prime}\left({{x^{*}}}\right)}}\right\|_{2}^{2}$
@@
$\leq$ $4{\mathbb{E}_{k}}\left[{\left\|{{r_{k}}-\nabla F\left({{x^{*}}}\right)}\right\|_{2}^{2}}\right]+4\left\|{{G^{\prime}\left({{x_{k+1}}}\right)}-{G^{\prime}\left({{x^{*}}}\right)}}\right\|_{2}^{2}$
$+16n{{\phi}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}+4n{{\phi}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}}$
@@
$+16n{{\phi}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{R}_{i}}}\right|}^{2}}}+4n{{\phi}^{2}}\sum\limits_{i\in\mathcal{R}}{{{\left|{{\mathcal{B}_{i}}}\right|}^{2}}},$
where the second inequality uses the results (48) and (50), and the last inequality is owing to Lemma 2 and Assumption 3. To proceed, recalling the definition of ${\partial G\left({{x_{k}}}\right)}$, it is not difficult to verify
$-\left\langle{{x_{k}}-{x^{*}},{G^{\prime}\left({{x_{k+1}}}\right)}-{G^{\prime}\left({{x^{*}}}\right)}}\right\rangle\leq 0,$ (68)
which is owing to the convexity of ${g}\left({\tilde{x}}\right)$, $\forall i\in\mathcal{R}$. Based on the relations (53) and (68), we know that
$-2{\mathbb{E}_{k}}\left[{\left\langle{{x_{k}}-{x^{*}},{r_{k}}+{G^{\prime}\left({{x_{k+1}}}\right)}+{\partial_{x}}\chi\left({{x_{k}}}\right)}\right\rangle}\right]$ (69)
@@
$-\!\!2{\mathbb{E}_{k}}\!\left[{\left\langle{{x_{k}}\!-\!{x^{*}},{G^{\prime}\left({{x_{k+1}}}\right)}\!-\!{G^{\prime}\left({{x^{*}}}\right)}\!+\!{\partial_{x}}\chi\left({{x_{k}}}\right)\!-\!{\partial_{x}}\chi\left({{x^{*}}}\right)}\right\rangle}\right]$